A system sitting on a floor was pushed with 100 N. The floor sliding coefficient of friction was 0.2. The system was a 25-kg cage that would push a 5-kg block in front with the 5-kg not touching the floor. The static friction coefficient between the cage and block was 0.5. Find the acceleration
Answer #1

From the FBD of small block we can write
F1 = ma
F1 = 5*a -----------(1)
f1 = mg (Since there is no motion in the vertical direction)
f1 = 5*9.81 = 49.05 N
Now the FBD of Cage
In vertical direction
R = Mg + f1 = (25*9.81)+49.05 = 294.3 N
Now in the horizontal direction
100 - F1 -f2 = Ma
where f2 = uk*R = 0.2*294.3 = 58.86 N
100 - 5a - 58.86 = 25a
30a = 41.14
a = 1.371 m/s2
5 kg 100 N